# FSFM: A Generalizable Face Security Foundation Model via Self-Supervised Facial Representation Learning

### v1

@@ E.2 Comparison with CLIP
Specifically, we borrow the CLIP image encoder, also a ViT-B/16 but pretrained on 400M (image, text) pairs, and finetune it on downstream face security tasks under the same settings as other base models. We supplement its results on deepfake detection, face anti-spoofing, and diffusion face forgery detection in Tab. 6, Tab. 7, and Tab. 8, respectively. We can observe that CLIP ViT-B transfers better than other base vision models on FAS and DiFF tasks, benefiting from the extensive scale of multi-modal supervision. However, directly applying CLIP ViT-B to DfD exhibits inferior generalization. In contrast, our proposed FSFM also consistently outperforms CLIP ViT-B across downstream face security tasks.
@@ F.1 Ablation studies
Effect of Model Scaling Tab. 9 also shows that FSFM benefits from larger model sizes when pretrained on FF++_o. The transfer performance on downstream face security tasks improves as the model scales up. Due to limited computing resources, we were unable to pretrain larger models on more face images, but we aspire to explore this in future work.
Effect of Loss Weight To explore the impact of the reconstruction and distillation losses, we vary various loss weights, i.e. and . Results in Fig. 9 show that the configuration (, ) performs better on challenging cross-dataset deepfake detection.

### v3

@@ E.3 Comparison with CLIP
Specifically, we borrow the CLIP image encoder, also a ViT-B but pretrained on the WIT dataset with 400M image& text pairs, and finetune it on downstream face security tasks under the same settings as other base models. We supplement the corresponding results on deepfake detection, face anti-spoofing, and diffusion face forgery detection in Tab. 7, Tab. 8, and Tab. 9, respectively. We can observe that CLIP ViT-B transfers better than other base vision models on FAS and DiFF tasks, benefiting from the extensive data scale of multi-modal supervision. However, directly applying CLIP ViT-B to DfD exhibits inferior generalization. In contrast, our proposed FSFM consistently outperforms CLIP ViT-B across downstream face security tasks.
@@ F.1 Ablation studies
Effect of Model Scaling Tab. 10 also shows that FSFM benefits from larger model sizes when pretrained on FF++_o. The transfer performance on downstream face security tasks improves as the model scales up. Due to limited computing resources, we were unable to pretrain larger models on more face images, e.g., pertaining ViT-L/16(303M) on the full VGGFace2 dataset, but we aspire to explore this in future work and update model zoos accordingly.
Effect of Loss Weight To explore the impact of the reconstruction and distillation losses, we vary various loss weights, i.e. and . Results in Fig. 9 show that the configuration (, ) performs better on challenging cross-dataset DfD.
@@ F.2 Visualizations
G Limitations
Despite the promising results demonstrated by FSFM across various face security tasks, our work has several limitations that warrant further exploration: Pretraining Dataset Bias FSFM is pretrained on large-scale facial images, and its performance can be affected by the quantity, diversity, and quality of the pretraining data. Pretraining on specific datasets like VGGFace2 [10] may inherit their biases (e.g., race, ethnicity, and age), potentially reducing fairness. Absence of Multi-modal Learning Since our work focuses on general visual face security, the current framework processes only image or frame data for downstream forgery&spoofing image&video detection, ignoring potential complementary signals (e.g., audio inconsistencies in deepfakes or physiological cues in spoofing), which could further enhance capabilities.